OpenStudy (anonymous):

find the range of f(x)=-sqrt(2x+4)+3

OpenStudy (mww):

We know that \[- \sqrt(2x+4) \] must always be negative or zero for the given domain. \[- \sqrt{2x+4} \le 0 \] Adding 3 to both sides of the inequality, we have: \[f(x) = 3 - \sqrt{2x+4} \le 3\] \[Range(f) = (-\infty, 3]\]